[Brahms-l] Shift report 20050210 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Feb 10 2005 - 16:03:37 EST
1. Beam is good in the morning, dumped at 12:00. No beam in the afternoon, "Machine Development" from 12:00-17:00

2. Setting Fs@4B3450 Mrs@45A1050
RunNo    Trig2   Trig3
13054    OFF      280k
13060    241k     116k    
13061    167k     150k
13062    117k     132k

TPC pedestal run 13065 
Non-TPC Pedestal run 13066

3. T2 got the problem, after Bob reset it, everything came back(see elog for info in detail)

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 10 16:04:06 2005

This archive was generated by hypermail 2.1.8 : Thu Feb 10 2005 - 16:04:21 EST